OpenStudy (anonymous):

Find the length of an arc on a circle of radius 5 corresponding to an angle of 135° give answer as fraction not a decimal.

OpenStudy (anonymous):

i am confused can you elaborate

OpenStudy (anonymous):

Which expression can be used to find the area of the circle below? A circle with radius labeled 9 centimeters A = (3.14)(2 • 9) A = (3.14)(2 • 3) A = (3.14)(32) A = (3.14)(92) can you answer this

OpenStudy (anonymous):

The arc length

OpenStudy (anonymous):

which it would be the last answer correct

OpenStudy (anonymous):

A= (3.14)(2*9)

OpenStudy (anonymous):

Which is the cost per square inch of a pizza with diameter 20 inches if the cost of the pizza is $16? Be sure to round your answer to the nearest cent. about 1 cent about 5 cents about twenty-five cents about eighty cents what about this i believe it is 25cents

OpenStudy (anonymous):

can you figure it out

OpenStudy (anonymous):

please

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

think